OpenStudy (anonymous):

Fabian has a savings account balance of $1,374.73. The interest rate on the account is 3.9% compounded twice a year. If he opened the account nine years ago, what was the value of his initial deposit?

OpenStudy (nurali):

A=P(1+rn)^nt 1374.73=P(1+0.0392)^2(9) Solve for P which is the initial amount.

OpenStudy (kropot72):

The following formula can be used to get the solution: \[A=P(1+\frac{r}{n})^{nt}\] where A is the amount after t years, P is the principal, r is the annual interest rate as a decimal and n is the number of compounding events per year. Plugging in values we get \[1374.73=P(1+\frac{0.039}{2})^{9\times 2}\] Now just solve for P.